027 10.0 points Suppose that 1.37 g of hydrogen is separated into electrons and protons, and also that the protons are placed at the Earths North Pole and the electrons are placed at the South Pole. What is the resulting compressional force on Earth? The value of the Coulomb constant is 8.98755 x 10 N m2/ and Avogadros number is 6.02214 x 10* mol-1. The radius of the Earth is 6.37 × 106 m. Answer in units of N.
Answer #1

Hydrogen molecule has a molar mass of 1g and has 2 electrons and protons. So total compressional force is

F_{compression}=N_{H{_2}}\times \frac{1}{4\pi\epsilon_0}\frac{2q^2}{r^2}\\ \implies F_{compression}=\frac{1.37}{1}\times 6.02214\times 10^{23}\times9\times10^9\times \frac{2\times (1.6\times 10^{-19})^2}{(2\times 6.37\times 10^6)^2}\\ \implies F_{compression}=2.34\times 10^{-18}N